A very large,thin flat plate of aluminium of area A has a total charge Q uniformly distributed over its surfaces. Assuming the same charge is spread uniformly over the upper surface of an otherwise identical glass plate, compare the electric fields just above the centre of the upper surface of each plate.
1
Expert's answer
2020-03-24T09:57:29-0400

The expression for the electric field on the surface of aluminum plate is,


"E_a=\frac{\sigma_a}{ \epsilon_0}=\frac{1}{ \epsilon_0}\frac{Q}{2A}"

The expression for the electric field on the surface of glass plate is,


"E_g=\frac{\sigma_g}{ \epsilon_0}=\frac{1}{ \epsilon_0}\frac{Q}{2A}"

"E_g=E_a"